### v1

# Existence of All Wilton Ripples of the Kawahara Equation

Abstract
The existence of all small-amplitude Wilton ripple solutions of the Kawahara equation is proven. These are periodic, traveling-wave solutions that bifurcate from a two-dimensional nullspace spanned by two distinct, co-propagating cosine waves. In contrast with previous results, the proof, which relies on a carefully constructed Lyapunov-Schmidt reduction, implies the existence of all small-amplitude Wilton ripples of the Kawahara equation, of which there are countably infinite. Though this result pertains only to the Kawahara equation, the method of proof likely extends to most nonlinear dispersive equations admitting Wilton ripple solutions.
Keywords: Kawahara Equation, Periodic Traveling Waves, Resonant Solutions, Wilton Ripples, Lyapunov-Schmidt Reduction, Nonsimple Local Bifurcations
Acknowledgments: This work was funded by the National Science Foundation through grant NSF-DMS-2402044.
1Division of Applied Mathematics, Brown University, RI, USA, ryan_creedon@brown.edu
1 Introduction
We consider the Kawahara (or super-KdV) equation
$u_{t}=\alpha u_{xxx}+\beta u_{5x}+\sigma\left(u^{2}\right)_{x},$ (1)
where $\alpha,\beta,$ and $\sigma$ are real, nonzero parameters. First derived in [13], the Kawahara equation models shallow water gravity-capillary waves near a critical value of the Bond number, a nondimensional quantity that compares gravity and surface tension. Owing to the competition between these forces, the Kawahara equation admits several interesting and physically-relevant solutions; see, for instance, [4]-[5] and [15, 21, 25]. Of importance in this work is a class of periodic, traveling-wave solutions called the Wilton ripples, which have a particularly rich history in the context of water waves [^1] More recently, these solutions have found practical relevance in the context of magnetohydrodynamics [26]. [24].
To motivate the mathematical behavior of Wilton ripples, we begin by searching for the periodic, traveling-wave solutions of the Kawahara equation. To this effect, we move to a traveling frame $x\rightarrow x-ct$ so that (1) becomes
$u_{t}=cu_{x}+\alpha u_{xxx}+\beta u_{5x}+\sigma\left(\sigma^{2}\right)_{x}.$ (2)
If $c$ is chosen to match the velocity of the traveling waves, then $u_{t}=0$, and we can integrate (2) once in $x$, yielding
$cu+\alpha u_{xx}+\beta u_{4x}+\sigma u^{2}=\mathcal{I},$ (3)
where $\mathcal{I}\in\mathbb{R}$ is a constant of integration. By exploiting a Galilean symmetry of (1), one can show that there exist boosts $u\rightarrow u+\xi_{1}$ and $c\rightarrow c+\xi_{2}$ of (3) such that $\mathcal{I}=0$ without loss of generality.
What remains is to study the $2\pi/\kappa$-periodic solutions of
$cu+\alpha u_{xx}+\beta u_{4x}+\sigma u^{2}=0,$ (4)
for all $\kappa>0$. In fact, by rescaling variables according to $x\rightarrow\kappa^{-1}x$, $u\rightarrow\alpha k^{2}\sigma^{-1}u$, $c\rightarrow\alpha k^{2}c$, and $\beta\rightarrow\alpha\kappa^{-2}\beta$, we can instead investigate the $2\pi$-periodic solutions of
$cu+u_{xx}+\beta u_{4x}+u^{2}=0,$ (5)
without loss of generality. Furthermore, because (5) is autonomous in $x$ and satisfies the discrete symmetry $x\rightarrow-x$, we need only consider its $2\pi$-periodic, even solutions.
If we further restrict these solutions to small-amplitude, (5) effectively linearizes to
$cu+u_{xx}+\beta u_{4x}=0.$ (6)
At first glance, it appears that the only [^2] It is also possible that $\cos(nx)$ solves (6) when $c=n^{2}-\beta n^{4}$ for any $n\in\mathbb{N}$. However, by folding $n$ into the definition of $\kappa$ before rescaling $x$, one can restrict to the $n=1$ case without loss of generality. real-valued solution of (6) with the requisite parity and periodicity occurs when $c=1-\beta$ and is given by $u=a\cos(x)$ for any $a\in\mathbb{R}$. However, if $\beta=1/(1+K^{2})$ for any $K\in\mathbb{N}\setminus\{1\}$, then $u=a\cos(x)+b\cos(Kx)$ satisfies (6) with $c=1-\beta$ for all $a,b\in\mathbb{R}$. Hence, depending on the value of $\beta$, we have two classes of small-amplitude, periodic traveling-wave solutions of the Kawahara equation:
Figure 1: (a) A Stokes wave of the Kawahara equation for $\beta=1/2$, (b) A Wilton ripple of the Kawahara equation for $\beta=1/5$, and (c) A Wilton ripple of the Kawahara equation for $\beta=1/10$. In all plots, the amplitude parameter $a=0.01$. The blue dots are numerically computed wave profiles using the continuation method presented in [23]. The red curves are the leading-order asymptotic behavior of the wave profiles according to Theorem 1 below. The wave profiles are normalized by $a$ for ease of comparisons.
  1. 1.
    When $\beta\neq 1/(1+K^{2})$, the periodic traveling-wave solutions bifurcate from a single cosine wave. For $|a|\ll 1$, these solutions take the form
    $u=a\cos(x)+u_{r}(x;a)\quad\textrm{and}\quad c=1-\beta+c_{r}(a),$ (7)
    where $u_{r}(x;a)$ is a smooth, $2\pi$-periodic function in $x$ that is real analytic in $a$ with $||u_{r}(x;a)||_{\textrm{L}^{2}}=\mathcal{O}\left(a^{2}\right)$ as $a\rightarrow 0$ and $c_{r}(a)$ is an even, real-analytic function in $a$ such that $|c_{r}(a)|=\mathcal{O}\left(a^{2}\right)$ as $a\rightarrow 0$. We call such solutions the Stokes waves of the Kawahara equation by analogy with the corresponding solutions of the surface gravity wave equations [22]. Their existence for sufficiently small $a$ was proven in [9] and their stability discussed in [7, 9, 17, 23]. The profile of a typical Stokes wave solution is provided in Figure 1(a).
  2. 2.
    When $\beta=1/(1+K^{2})$, the periodic traveling-wave solutions bifurcate from a superposition of two co-propagating cosine waves. These are the Wilton ripples of the Kawahara equation, named by analogy with the corresponding solutions of the gravity-capillary wave equations [24]. Unlike the Stokes waves, Wilton ripples arise from a two-dimensional bifurcation problem, and as a consequence, their expansion for small $a$ is different than that of the Stokes waves. Indeed, based on the formal asymptotic expansions derived in [10, 14], the Wilton ripple solutions of the Kawahara equation take the form
    $u=a\Big{(}\cos(x)+b(a)\cos(Kx)\Big{)}+u_{r}(x;a)\quad\textrm{and}\quad c=1-\beta+c_{r}(a),$ (8)
    where $b(a)$ is a real analytic function of $a$, $u_{r}(x;a)$ is a smooth, $2\pi$-periodic function of $x$ that is real analytic in $a$, and $c_{r}(a)$ is a real analytic function of $a$. The leading-order behaviors of $b(a)$, $u_{r}(x;a)$, and $c_{r}(a)$ as $a\rightarrow 0$ depend on $K$ according to Theorem 1 below. Profiles of typical Wilton ripple solutions are provided in Figures 1(b) and 1(c).
Partial existence proofs have emerged for Wilton ripples of the Kawahara equation and other nonlinear dispersive equations [2]-[3],[8], including those for gravity-capillary water waves [12, 16],[19]-[20]. Nearly all of these proofs, however, restrict to the simplest possible ripple that bifurcates from a superposition of $\cos(x)$ and $\cos(2x)$. In the context of the Kawahara equation, this is the special case when $K=2$.
A notable exception is [8], which adopts a new framework for studying Wilton ripples of a gravity-capillary Whitham equation. In this framework, both $a$ and $b$ are fixed, and the remaining parameters in the governing equation of the Wilton ripples [^3] This would be $\beta$ in the Kawahara equation. become functions of $a$ and $b$. As a consequence, one can prove the existence of a two-dimensional “sheet” of Wilton ripples parameterized by $a$ and $b$. Recent numerical and formal asymptotic investigations by [1] have applied this new framework to the Kawahara equation, and the results suggest that the Wilton ripples discussed herein are embedded in this sheet of solutions (at least for $K=2$ and $K=3$). However, no rigorous proof of this embedding has been established. Thus, the existence of Wilton ripples for all $K\in\mathbb{N}\setminus\{1\}$ remains an open problem. It is the objective of this work to prove this fact. To the author’s knowledge, this is the first proof of its kind to demonstrate the existence of all Wilton ripple solutions of a nonlinear dispersive equation.
In contrast to many previous existence proofs, which rely on Banach fixed point arguments or the method of majorants, this proof relies on a carefully constructed Lyapunov-Schmidt reduction to prove that each Wilton ripple depends analytically on a small-amplitude parameter. It is anticipated that this part of the proof extends with minimal changes to Wilton ripples of any nonlinear dispersive equation, including the gravity-capillary water wave equations. However, to finish the existence proof, one must show that the function $b(a)$ in (8) does not vanish identically for each $K\in\mathbb{N}\setminus\{1\}$. In order to show this, we need high-order asymptotic expansions of all Wilton ripples. For general nonlinear dispersive equations, these expansions have not been carried out, but recent work by [14] has derived these expansions for the Kawahara equation. Using these expansions, we can prove our main result:
Theorem 1.
For each $K\in\mathbb{N}\setminus\{1\}$, the Kawahara equation has the following Wilton ripple solutions:
  • If $K=2$ (so that $\beta=1/5$), there exist two $a$-parameter families of solutions $u_{+}(x;a)$ and $u_{-}(x;a)$ that solve (5) for sufficiently small $a\in\mathbb{R}$. The solutions have velocity
    $c_{\pm}=\frac{4}{5}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a),$ (9)
    respectively, where $\tilde{c_{r}}^{\pm}(a)$ are real analytic functions such that $\tilde{c_{r}}^{\pm}(0)=\pm 1.$ The solutions take the form
    $u_{\pm}(x;a)=a\Big{(}\cos(x)+\frac{1}{\sqrt{2}}\tilde{b}^{\pm}(a)\cos(2x)\Big{)}+u_{r}\Big{(}x;a,\frac{a}{\sqrt{2}}\tilde{b}^{\pm}(a),-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a)\Big{)},$ (10)
    where $\tilde{b}^{\pm}(a)$ are real analytic functions such that $\tilde{b}^{\pm}(0)=\pm 1$ and $u_{r}\in\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$ is a real analytic function in its final three arguments.
  • If $K=3$ (so that $\beta=1/10$), there exist three a-parameter families of solutions $u_{\sigma}(x;a)$ for $\sigma\in\{1,2,3\}$ that solve (5) for sufficiently small $a\in\mathbb{R}$. The solutions have velocity
    $c_{\sigma}=\frac{9}{10}+a^{2}\tilde{c_{r}}^{(\sigma)}(a),$ (11)
    where $\tilde{c_{r}}^{(\sigma)}(a)$ are real analytic functions such that
    $\tilde{c_{r}}^{(1)}(0)\approx 4.27863,\quad\tilde{c_{r}}^{(2)}(0)\approx 1.48289,\quad\textrm{and}\quad\tilde{c_{r}}^{(3)}(0)\approx 0.37396.$ (12)
    The solutions take the form
    $u_{\sigma}(x;a)=a\Big{(}\cos(x)+\tilde{b}^{(\sigma)}(a)\cos(3x)\Big{)}+u_{r}\Big{(}x;a,a\tilde{b}^{(\sigma)}(a),a^{2}\tilde{c_{r}}^{(\sigma)}(a)\Big{)},$ (13)
    where $\tilde{b}^{(\sigma)}(a)$ are real analytic functions such that
    $\tilde{b}^{(1)}(0)\approx-1.78374,\quad\tilde{b}^{(2)}(0)\approx-0.54488,\quad\textrm{and}\quad\tilde{b}^{(3)}(0)\approx 0.59468,$ (14)
    and $u_{r}\in\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$ is a real analytic function in its final three arguments.
  • If $K\geq 4$ (so that $\beta=1/(1+K^{2})$), there exists one $a$-parameter family of solutions $u(x;a)$ that solves (5) for sufficiently small $a\in\mathbb{R}$. The solutions have velocity
    $c=\frac{K^{2}}{K^{2}+1}+a^{2}\tilde{c_{r}}(a),$ (15)
    where $\tilde{c_{r}}(a)$ is a real analytic function such that
    $\tilde{c_{r}}(0)=\frac{(K^{2}+1)(5K^{2}-24)}{6K^{2}(K^{2}-4)}.$ (16)
    The solutions take the form
    $u(x;a)=a\Big{(}\cos(x)+\tilde{b}(a)\cos(Kx)\Big{)}+u_{r}\Big{(}x;a,a\tilde{b}(a),a^{2}\tilde{c_{r}}(a)\Big{)},$ (17)
    where $\tilde{b}(a)$ is a nonzero, real analytic function such that $\tilde{b}(a)=\mathcal{O}\left(a^{K-3}\right)$ as $a\rightarrow 0$ and $u_{r}\in\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$ is a real analytic function in its final three arguments.
2 Proof of Theorem 1
2.1 Notation and Set-Up
Fix $K\in\mathbb{N}\setminus\{1\}$ and let $\beta=1/(1+K^{2})$. On the interval $\mathbb{T}=[-\pi,\pi]$, define the Hilbert space
$\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T}):=\left\{f:f(-\pi)=f(\pi),~{}f(-x)=f(x),~{}\textrm{and}~{}\sum_{k=0}^{\infty}(1+k^{2})^{4}\left|\widehat{f}_{k}\right|^{2}<+\infty\right\},$ (18)
where
[…]
are the Fourier cosine coefficients of $f$. Define the nonlinear operator $F:\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})\times\mathbb{R}\rightarrow\textrm{L}^{2}_{\textrm{per},\textrm{even}}(\mathbb{T})$ according to
$F(u,c)=(c+\mathcal{L})u+u^{2},\quad\textrm{where}\quad\mathcal{L}:=\partial_{x}^{2}+\beta\partial_{x}^{4}.$ (20)
Then, (5) becomes the operator equation
$F(u,c)=0.$ (21)
Over the given domain $\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$, the linear operator $\mathcal{L}$ is bounded for all $c\in\mathbb{R}$ and is self-adjoint with respect to the standard $\textrm{L}^{2}(\mathbb{T})$ inner-product:
[…]
As discussed in the Introduction, $c_{0}+\mathcal{L}$ has a two-dimensional nullspace spanned by $\cos(x)$ and $\cos(Kx)$ when $c_{0}:=1-\beta$. Owing to self-adjointness, the cokernel of $c_{0}+\mathcal{L}$ is also spanned by $\cos(x)$ and $\cos(Kx)$, which guarantees that $c_{0}+\mathcal{L}$ is a Fredholm operator of index zero.
In what follows, we decompose $c$ according to
$c=c_{0}+c_{r},$ (23)
where $c_{r}\in\mathbb{R}$ is conceived as a small correction to $c_{0}$. We also decompose $u\in\textrm{H}^{4}_{\textrm{per,even}}(\mathbb{T})$ by
$u=u_{0}+u_{r},$ (24)
where $u_{0}\in\textrm{Null}\{c_{0}+\mathcal{L}\}$ and $u_{r}\in\textrm{Null}\{c_{0}+\mathcal{L}_{0}\}^{\perp}$. (Here, orthogonality is taken with respect to the inner-product (22).) By design,
$u_{0}(x)=a\cos(x)+b\cos(Kx)$ (25)
for some $a,b\in\mathbb{R}$, while $u_{r}$ is treated as a small correction to $u_{0}$ involving Fourier cosines modes that are complementary to $\cos(x)$ and $\cos(Kx)$.
Substituting (23) and (24) into (20) yields a new operator equation
$G(u_{r},a,b,c_{r})=0,$ (26)
where $G:\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\times\mathbb{R}^{3}\rightarrow\textrm{L}^{2}_{\textrm{per},\textrm{even}}(\mathbb{T})$ such that
$G(u_{r},a,b,c_{r})$ $:=F\left(a\cos(x)+b\cos(Kx)+u_{r},c_{0}+c_{r}\right)$
$~{}=(c_{0}+\mathcal{L})u_{r}+c_{r}\left(a\cos(x)+b\cos(Kx)+u_{r}\right)+\left(a\cos(x)+b\cos(Kx)+u_{r}\right)^{2}.$ (27)
We are now ready to proceed with Lyapunov-Schmidt reduction on (26).
2.2 Solving the Auxiliary Equation
We introduce the projection operators $\mathcal{P}$ and $\mathcal{Q}$ according to
$\mathcal{P}f$ $=\widehat{f}_{1}\cos(x)+\widehat{f}_{K}\cos(Kx),$ (28a)
$\mathcal{Q}f$ $=f-\mathcal{P}f,$ (28b)
which are valid for all $f\in\textrm{L}^{2}_{\textrm{per,even}}(\mathbb{T})$. Here, the operator $\mathcal{P}$ projects any function in $\textrm{L}^{2}_{\textrm{per,even}}(\mathbb{T})$ onto the cokernel of $c_{0}+\mathcal{L}$, while $\mathcal{Q}$ projects onto the range of $c_{0}+\mathcal{L}$.
Applying $\mathcal{Q}$ to (26) yields a new operator equation
$\mathcal{A}(u_{r},a,b,c_{r})=0,$ (29)
where
$\mathcal{A}(u_{r},a,b,c_{r}):=\mathcal{Q}G(u_{r},a,b,c_{r})=c_{r}u_{r}+\mathcal{Q}(c_{0}+\mathcal{L})u_{r}+\mathcal{Q}\left(a\cos(x)+b\cos(Kx)+u_{r}\right)^{2}.$ (30)
We call (29) the auxiliary equation. From (30), it follows that
$\mathcal{A}(0,0,0,0)=0.$ (31)
Moreover, since the operator $F$ in (20) is real analytic with respect to $u$ and $c$ and the projections $\mathcal{P}$ and $\mathcal{Q}$ are bounded linear operators, $\mathcal{A}$ is real analytic with respect to $u_{r}$, $a$, $b$, and $c_{r}$. A straightforward calculation shows that
$\partial_{u_{r}}\mathcal{A}(0,0,0)=\mathcal{Q}(c_{0}+\mathcal{L}),$ (32)
where $\partial_{u_{r}}$ denotes the Fréchet derivative with respect to $u_{r}$. The resulting operator $\mathcal{Q}(c_{0}+\mathcal{L}):\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\rightarrow\textrm{Range}\{c_{0}+\mathcal{L}\}$ has bounded inverse
$\left(\mathcal{Q}(c_{0}+\mathcal{L})\right)^{-1}(f(x))=\sum_{k\in\mathbb{N}_{0}\setminus\{1,K\}}\left(\frac{\widehat{f}_{k}}{c_{0}-k^{2}+\beta k^{4}}\right)\cos(kx).$ (33)
As a consequence, the implicit function theorem gives the existence and uniqueness of
$u_{r}=u_{r}(x;a,b,c_{r}),$ (34)
such that, for sufficiently small $a$, $b$, and $c_{r}$, we have (i) $u_{r}(x;a,b,c_{r})\in\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\subset\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$, (ii) $\mathcal{A}(u_{r}(x;a,b,c_{r}),a,b,c_{r})=0$, (iii) $u_{r}(x;0,0,0)=0$, and (iv) $u_{r}(x;a,b,c_{r})$ is jointly real analytic with respect to $a$, $b$, and $c_{r}$.
By analyticity, (34) admits a Taylor series expansion of the form
$u_{r}(x;a,b,c_{r})$ $=u_{r}^{(1,0,0)}(x)a+u_{r}^{(0,1,0)}(x)b+u_{r}^{(0,0,1)}(x)c_{r}+u_{r}^{(2,0,0)}(x)a^{2}+u_{r}^{(1,1,0)}(x)ab$
$\quad+u_{r}^{(0,2,0)}(x)b^{2}+u_{r}^{(0,1,1)}(x)bc_{r}+u_{r}^{(0,0,2)}(x)c_{r}^{2}+u_{r}^{(1,0,1)}(x)ac_{r}+R_{3}(x;a,b,c_{r}),$ (35)
where $u_{r}^{(i,j,k)}(x)$ are explicitly computed in the Appendix and $R_{3}(x;a,b,c_{r})\in\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\subset\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$ such that
$\Big{|}\Big{|}R_{3}(x;a,b,c_{r})\Big{|}\Big{|}_{\textrm{H}^{4}(\mathbb{T})}=\mathcal{O}\left((|a|+|b|+|c_{r}|)^{3}\right)\quad\textrm{as}\quad a,b,c_{r}\rightarrow 0;$ (36)
see [6] for details.
2.3 Solving the Bifurcation Equation
Applying the projection $\mathcal{P}$ to (26) yields a new operator equation
$\mathcal{B}(u_{r},a,b,c_{r})=0,$ (37)
where $\mathcal{B}:\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\crossproduct\mathbb{R}^{3}\rightarrow\textrm{Coker}\{c_{0}+\mathcal{L}\}$ such that
$\mathcal{B}(u_{r},a,b,c_{r})$ $:=\mathcal{P}G(u_{r},a,b,c_{r})$
$~{}=c_{r}\left(a\cos(x)+b\cos(Kx)\right)+\mathcal{P}\left(a\cos(x)+b\cos(Kx)+u_{r}\right)^{2}.$ (38)
Substituting $u_{r}=u_{r}(x;a,b,c_{r})$ into (37) and separating terms proportional to $\cos(x)$ and $\cos(Kx)$, we arrive at a finite-dimensional operator equation
~$\tilde{\mathcal{B}}(a,b,c_{r})=0,$ (39)
where $\tilde{B}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$ such that
~$\tilde{\mathcal{B}}(a,b,c_{r})=\begin{pmatrix}\mathcal{F}_{1}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]+ac_{r}\\ \mathcal{F}_{K}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]+bc_{r}\end{pmatrix}.$ (40)
We call (39) the bifurcation equation.
Owing to the analyticity of $u_{r}(x;a,b,c_{r})$, we expand $\tilde{\mathcal{B}}$ for sufficiently small $a$, $b$, and $c_{r}$. To this end, we let $\omega\in\{1,K\}$ and consider
$\mathcal{F}_{\omega}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]=\mathcal{T}_{\omega,1}+\mathcal{T}_{\omega,2}+\mathcal{T}_{\omega,3},$ (41)
where
$\mathcal{T}_{\omega,1}$ $=\mathcal{F}_{\omega}\left[\left(a\cos(x)+b\cos(Kx)\right)^{2}\right],$ (42a)
$\mathcal{T}_{\omega,2}$ $=\mathcal{F}_{\omega}\left[2\left(a\cos(x)+b\cos(Kx)\right)u_{r}(x;a,b,c_{r})\right],$ (42b)
$\mathcal{T}_{\omega,3}$ $=\mathcal{F}_{\omega}\left[u_{r}(x;a,b,c_{r})^{2}\right].$ (42c)
For sufficiently small $a$, $b$, and $c_{r}$, we have
$\left|\mathcal{T}_{\omega,3}\right|\leq||u_{r}(x;a,b,c_{r})||^{2}_{\textrm{L}^{2}(\mathbb{T})}\leq||u_{r}(x;a,b,c_{r})||^{2}_{\textrm{H}^{4}(\mathbb{T})}.$ (43)
Moreover, because the linear terms in the Taylor series of $u_{r}(x;a,b,c_{r})$ vanish for all $K\in\mathbb{N}\setminus\{1\}$ by (88) and (91) in the Appendix, we must have
$||u_{r}(x;a,b,c_{r})||_{\textrm{H}^{4}(\mathbb{T})}\leq C\left(|a|+|b|+|c_{r}|\right)^{2}$ (44)
for some constant $C=C(K)>0$ by Taylor’s theorem. Thus,
$\mathcal{T}_{\omega,3}=\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\quad\textrm{as}\quad a,b,c_{r}\rightarrow 0,$ (45)
regardless of $\omega$. In contrast, $\mathcal{T}_{\omega,1}$ and $\mathcal{T}_{\omega,2}$ contribute terms lower than fourth degree in $a$, $b$, and $c_{r}$ when expanded. The coefficients of these terms are different depending on whether $K=2$, $K=3$, or $K\geq 4$. Thus, we must expand the bifurcation equation in each of these cases separately.
2.3.1 Case 1: $K=2$
A direct calculation of (42a) when $K=2$ yields
$\mathcal{T}_{\omega,1}=\begin{cases}ab&\omega=1\\ \frac{1}{2}a^{2}&\omega=2\end{cases}.$ (46)
Turning to $\mathcal{T}_{\omega,2}$, we substitute the expansion (35) into (42b), giving us
$\mathcal{T}_{\omega,2}$ $=\mathcal{F}_{\omega}\left[2\big{(}a\cos(x)+b\cos(2x)\big{)}\big{(}u^{(2,0,0)}(x)a^{2}+u^{(1,1,0)}(x)ab+u^{(0,2,0)}(x)b^{2}\big{)}\right]+\mathcal{R}_{\omega,2},$ (47)
where $u^{(i,j,k)}(x)$ are given in (90a)-(90c) and
$\mathcal{R}_{\omega,2}=\mathcal{F}_{\omega}\left[2\big{(}a\cos(x)+b\cos(2x)\big{)}R_{3}(x;a,b,c_{r})\right].$ (48)
Using a combination of the Cauchy-Schwarz inequality and Taylor’s theorem, we find
$|\mathcal{R}_{\omega,2}|$ $\leq 2\big{(}|a|+|b|\big{)}||R_{3}(x;a,b,c_{r})||_{\textrm{L}^{1}(\mathbb{T})},$
$\leq 2\sqrt{2\pi}(|a|+|b|)||R_{3}(x;a,b,c_{r})||_{\textrm{L}^{2}(\mathbb{T})},$
$\leq 2\sqrt{2\pi}(|a|+|b|)||R_{3}(x;a,b,c_{r})||_{\textrm{H}^{4}(\mathbb{T})},$
$\leq 2\sqrt{2\pi}C(|a|+|b|+|c_{r}|)^{3}(|a|+|b|),$
$\leq 2\sqrt{2\pi}C(|a|+|b|+|c_{r}|)^{4},$ (49)
which is valid for sufficiently small $a$, $b$, and $c_{r}$. It follows that
$\mathcal{R}_{\omega,2}=\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\quad\textrm{as}\quad a,b,c_{r}\rightarrow 0.$ (50)
The remaining term in (47) can be calculated explicitly with the aid of Mathematica [^4] See the Mathematica companion file CompanionToKawaharaWiltonRipples.nb for details.. We find
$\mathcal{T}_{\omega,2}=\begin{cases}-\frac{5}{4}a^{3}-\frac{11}{8}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}&\omega=1\\ -\frac{11}{8}a^{2}b-\frac{91}{72}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}&\omega=2\end{cases},$ (51)
where we have used $\beta=1/(1+K^{2})=1/5$ and $c_{0}=1-\beta=4/5$ to simplify. Substituting (45), (46), and (51) into (39) yields the following expansion of the bifurcation equation:
~$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ab+ac_{r}-\frac{5}{4}a^{3}-\frac{11}{8}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ \frac{1}{2}a^{2}+bc_{r}-\frac{11}{8}a^{2}b-\frac{91}{72}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (52)
Comparison of the dominant terms in (52) suggests the following rescaling:
$b=\frac{a}{\sqrt{2}}\tilde{b},\quad c_{r}=-\frac{a}{\sqrt{2}}\tilde{c_{r}},\quad\textrm{and}\quad\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\frac{1}{a^{2}}\tilde{\mathcal{B}}\left(a,\frac{a}{\sqrt{2}}\tilde{b},-\frac{a}{\sqrt{2}}\tilde{c_{r}}\right).$ (53)
Then, (52) becomes
$\tilde{\tilde{B}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}\frac{1}{\sqrt{2}}(\tilde{b}-\tilde{c_{r}})-\frac{5}{4}a-\frac{11}{16}a\tilde{b}^{2}+\mathcal{O}\left(|a|^{2}\left(1+\frac{1}{\sqrt{2}}|\tilde{b}|+\frac{1}{\sqrt{2}}|\tilde{c_{r}}|\right)^{4}\right)\\ \frac{1}{2}(1-\tilde{b}\tilde{c_{r}})-\frac{11}{8\sqrt{2}}a\tilde{b}-\frac{91}{144\sqrt{2}}a\tilde{b}^{3}+\mathcal{O}\left(|a|^{2}\left(1+\frac{1}{\sqrt{2}}|\tilde{b}|+\frac{1}{\sqrt{2}}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (54)
From (54), we have that ~~$\tilde{\tilde{\mathcal{B}}}(0,\pm 1,\pm 1)=0$. The corresponding Jacobian evaluated at $(0,\pm 1,\pm 1)$ is
~~$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,\pm 1,\pm 1)=\begin{pmatrix}\frac{1}{\sqrt{2}}&-\frac{1}{\sqrt{2}}\\ \mp 1&\mp 1\end{pmatrix},$ (55)
which is invertible in both cases. The implicit function theorem guarantees the existence and uniqueness of
$\tilde{b}^{\pm}=\tilde{b}^{\pm}(a)\quad\textrm{and}\quad\tilde{c_{r}}^{\pm}=\tilde{c_{r}}^{\pm}(a),$ (56)
such that, for sufficiently small $a$, we have (i) ~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}^{\pm}(a),\tilde{c_{r}}^{\pm}(a))=0$, (ii) $\tilde{b}^{\pm}(0)=\pm 1$ and $\tilde{c_{r}}^{\pm}(0)=\pm 1$, and (iii) $\tilde{b}^{\pm}(a)$ and $\tilde{c_{r}}^{\pm}(a)$ are real analytic in $a$.
Having solved both the auxiliary and bifurcation equations when $K=2$, we conclude the existence of sufficiently small-amplitude Wilton ripple solutions of (5) in this case. In particular, if $K=2$ and $\beta=1/(1+K^{2})=1/5$, there exist two unique $a$-parameter family of solutions to (5) for sufficiently small $a\in\mathbb{R}$. These solutions travel at velocity
$c=c_{0}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a)=\frac{4}{5}-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a),$ (57)
respectively, and take the form
$u_{\pm}(x;a)=a\Big{(}\cos(x)+\frac{1}{\sqrt{2}}\tilde{b}^{\pm}(a)\cos(2x)\Big{)}+u_{r}\Big{(}x;a,\frac{a}{\sqrt{2}}\tilde{b}^{\pm}(a),-\frac{a}{\sqrt{2}}\tilde{c_{r}}^{\pm}(a)\Big{)},$ (58)
where $u_{r}\in\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$ is real analytic in its parameters with Taylor expansion (35) and coefficients (90a)-(90c). The functions $\tilde{b}^{\pm}(a)$ and $\tilde{c_{r}}^{\pm}(a)$ are real analytic in $a$ such that $\tilde{b}^{\pm}(0)=\pm 1$ and $\tilde{c_{r}}(0)=\pm 1$. Consequently, $u_{\pm}(x;a)$ are real analytic in $a$ with values in $\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$. In fact, by rewriting (5) as
$\partial_{x}^{4}u_{\pm}(x;a)=-\beta^{-1}\left(cu_{\pm}(x;a)+\partial_{x}^{2}u_{\pm}(x;a)+u_{\pm}(x;a)^{2}\right),$ (59)
we infer $\partial_{x}^{4}u_{\pm}(x;a)\in\textrm{H}^{2}_{\textrm{per},\textrm{even}}(\mathbb{T})$ and, hence, $u_{\pm}(x;a)\in\textrm{H}^{6}_{\textrm{per},\textrm{even}}(\mathbb{T})$. Repeating this argument indefinitely, one concludes $u_{\pm}(x;a)\in\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$, completing the proof of Theorem 1 when $K=2$.
2.3.2 Case 2: $K=3$
When $K=3$, a direct calculation shows
$\mathcal{T}_{\omega,1}=0,$ (60)
regardless of $\omega$. Turning to $\mathcal{T}_{\omega,2}$, we substitute (92a)-(92c) into (47) with $K=3$ to obtain the following expansion:
$\mathcal{T}_{\omega,2}=\begin{cases}-\frac{7}{9}a^{3}+a^{2}b-\frac{34}{63}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}&\omega=1\\ \frac{1}{3}a^{3}-\frac{34}{63}a^{2}b-\frac{211}{189}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}&\omega=3\end{cases},$ (61)
where we have used $\beta=1/(1+K^{2})=1/10$ and $c_{0}=1-\beta=9/10$ to simplify the coefficients above. The remainder terms are estimated exactly as in the previous case.
Given (61), the bifurcation equation expands as
~$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ac_{r}-\frac{7}{9}a^{3}+a^{2}b-\frac{34}{63}ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ bc_{r}+\frac{1}{3}a^{3}-\frac{34}{63}a^{2}b-\frac{211}{189}b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (62)
Balancing the quadratic and cubic terms above suggests the following rescaling:
$b=a\tilde{b},\quad c_{r}=a^{2}\tilde{c_{r}},\quad\textrm{and}\quad\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\frac{1}{a^{3}}\tilde{B}(a,a\tilde{b},a^{2}\tilde{c_{r}}),$ (63)
leading us to a new bifurcation equation
~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}-\frac{7}{9}+\tilde{c_{r}}+\tilde{b}-\frac{34}{63}\tilde{b}^{2}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\\ \frac{1}{3}-\frac{34}{63}\tilde{b}+\tilde{b}\tilde{c_{r}}-\frac{211}{189}\tilde{b}^{3}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (64)
When $a=0$, we obtain a solution to the above provided
$\tilde{c_{r}}=\frac{7}{9}-\tilde{b}+\frac{34}{63}\tilde{b}^{2}\quad\textrm{and}\quad\tilde{b}^{3}+\tilde{b}^{2}-\frac{5}{21}\tilde{b}-\frac{1}{3}=0.$ (65)
Since the discriminant of the $\tilde{b}$ equation is positive, we anticipate three distinct, real solutions. Arranged from least to greatest, we denote these solutions by $\tilde{b}^{(\sigma)}_{0}$ for $\sigma=1,~{}2,~{}\textrm{and}~{}3$, respectively. Numerical methods reveal
$\tilde{b}^{(1)}_{0}\approx-1.78374,\quad\tilde{b}^{(2)}_{0}\approx-0.54488,\quad\textrm{and}\quad\tilde{b}^{(3)}_{0}\approx 0.59468.$ (66)
For each $\sigma$, denote $\tilde{c_{r,0}}^{(\sigma)}$ by the evaluation of (65) at $\tilde{b}_{0}^{(\sigma)}$. Using the numerical values (66), we find
$\tilde{c_{r,0}}^{(1)}\approx 4.27863,\quad\tilde{c_{r,0}}^{(2)}\approx 1.48289,\quad\textrm{and}\quad\tilde{c_{r,0}}^{(3)}\approx 0.37396.$ (67)
By construction, it follows necessarily that
~~$\tilde{\tilde{\mathcal{B}}}\left(0,\tilde{b}_{0}^{(\sigma)},\tilde{c_{r,0}}^{(\sigma)}\right)=0,$ (68)
for each $\sigma\in\{1,2,3\}$. The corresponding Jacobians are
~~$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,\tilde{b}_{0}^{(\sigma)},\tilde{c_{r,0}}^{(\sigma)})=\begin{pmatrix}1&1\\ -\frac{34}{63}&\tilde{b}_{0}^{(\sigma)}\end{pmatrix},$ (69)
which are invertible since $\tilde{b}_{0}^{(\sigma)}\neq-34/63$ for all $\sigma$ by contradiction. The implicit function theorem then gives the existence and uniqueness of
$\tilde{b}^{(\sigma)}=\tilde{b}^{(\sigma)}(a)\quad\textrm{and}\quad\tilde{c_{r}}^{(\sigma)}=\tilde{c_{r}}^{(\sigma)}(a),$ (70)
for each $\sigma\in\{1,2,3\}$ and sufficiently small $a$ such that (i) ~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}^{(\sigma)}(a),\tilde{c_{r}}^{(\sigma)}(a))=0$, (ii) $\tilde{b}^{(\sigma)}(0)=\tilde{b}_{0}^{(\sigma)}$ and $\tilde{c_{r}}^{(\sigma)}(0)=\tilde{c_{r,0}}^{(\sigma)}$, and (iii) $\tilde{b}^{(\sigma)}(a)$ and $\tilde{c_{r}}^{(\sigma)}(a)$ are real analytic in $a$.
We conclude the existence of sufficiently small-amplitude Wilton ripples of (5) when $K=3$, i.e., $\beta=1/10$. In this case, there exist three unique $a$-parameter family of solutions to (5) for sufficiently small $a\in\mathbb{R}$. These solutions travel at velocity
$c=c_{0}+a^{2}\tilde{c_{r}}^{(\sigma)}(a)=\frac{9}{10}+a^{2}\tilde{c_{r}}^{(\sigma)}(a),$ (71)
for $\sigma\in\{1,2,3\}$, and their corresponding profiles are
$u_{\sigma}(x;a)=a\Big{(}\cos(x)+\tilde{b}^{(\sigma)}(a)\cos(3x)\Big{)}+u_{r}\Big{(}x;a,a\tilde{b}^{(\sigma)}(a),a^{2}\tilde{c_{r}}^{(\sigma)}(a)\Big{)},$ (72)
where $u_{r}\in\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$ is real analytic in its parameters with Taylor expansion (35) and coefficients (92a)-(92c) evaluated at $K=3$. The functions $\tilde{b}^{(\sigma)}(a)$ and $\tilde{c_{r}}^{(\sigma)}(a)$ are real analytic for sufficiently small $a$ and satisfy
$\tilde{b}^{(\sigma)}(0)=\tilde{b}_{0}^{(\sigma)}\quad\textrm{and}\quad\tilde{c_{r}}^{(\sigma)}(0)=\tilde{c_{r,0}}^{(\sigma)}.$ (73)
By the same bootstrap argument as before, we conclude $u_{\sigma}(x;a)$ are real analytic functions of $a$ with values in $\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$ for each $\sigma$, completing the proof of Theorem 1 when $K=3$.
2.3.3 Case 3: $K\geq 4$
As in the $K=3$ case, we find
$\mathcal{T}_{\omega,1}=0,$ (74)
regardless of $\omega$. Substituting (92a)-(92c) into (47) yields the following expansion for $K\geq 4$:
$\mathcal{T}_{\omega,2}=\begin{cases}v_{3,0,0}(K)a^{3}+v_{1,2,0}(K)ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}&\omega=1\\ w_{2,1,0}(K)a^{2}b+w_{0,3,0}(K)b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}&\omega=K\end{cases},$ (75)
where
$v_{3,0,0}(K)$ $=-\frac{(K^{2}+1)(5K^{2}-24)}{6K^{2}(K^{2}-4)},$ (76a)
$v_{1,2,0}(K)=w_{2,1,0}(K)$ $=-\frac{(K^{2}+1)(4K^{4}-27K^{2}+4)}{K^{2}(K^{2}-4)(4K^{2}-1)},$ (76b)
$w_{0,3,0}(K)$ $=-\frac{(K^{2}+1)(24K^{2}-5)}{6K^{2}(4K^{2}-1)}.$ (76c)
Given (75), the bifurcation equation expands as
~$\tilde{\mathcal{B}}(a,b,c_{r})$ $=\begin{pmatrix}ac_{r}+v_{3,0,0}(K)a^{3}+v_{1,2,0}(K)ab^{2}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\\ bc_{r}+w_{2,1,0}(K)a^{2}b+w_{0,3,0}(K)b^{3}+\mathcal{O}\left(|a|+|b|+|c_{r}|\right)^{4}\end{pmatrix}=0.$ (77)
We rescale variables as in (63) and arrive at the new bifurcation equation
~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b},\tilde{c_{r}})=\begin{pmatrix}v_{3,0,0}(K)+\tilde{c_{r}}+v_{1,2,0}(K)\tilde{b}^{2}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\\ w_{2,1,0}(K)\tilde{b}+\tilde{c_{r}}\tilde{b}+w_{0,3,0}(K)\tilde{b}^{3}+\mathcal{O}\left(|a|\left(1+|\tilde{b}|+|a|^{2}|\tilde{c_{r}}|\right)^{4}\right)\end{pmatrix}=0.$ (78)
When $a=0$, we obtain a solution to the bifurcation equation provided
$\tilde{c_{r}}=-v_{3,0,0}(K)-v_{1,2,0}(K)\tilde{b}^{2},\quad\quad\quad$ (79a)
$\big{(}w_{0,3,0}(K)-v_{1,2,0}(K)\big{)}\tilde{b}^{3}-\big{(}v_{3,0,0}(K)-w_{2,1,0}(K)\big{)}\tilde{b}$ $=0.$ (79b)
Notably, the cubic for $\tilde{b}$ is simpler in the present case than in the $K=3$ case. This is because the leading-order Fourier modes of $u_{r}(x;a,b,c_{r})$ in $\mathcal{T}_{\omega,2}$ have more resonant interactions with $\cos(x)$ and $\cos(Kx)$ when $K=3$. The obvious solution of the cubic is $\tilde{b}=0$. The other solutions satisfy
$\tilde{b}^{2}=\frac{v_{3,0,0}(K)-w_{2,1,0}(K)}{w_{0,3,0}(K)-v_{1,2,0}(K)}=-\frac{K^{2}(4K^{2}-61)}{61K^{2}-4}.$ (80)
For $K\geq 4$, the right-hand side of this equation is always negative. Thus, the remaining two solutions of the cubic are imaginary and will not contribute real-valued Wilton ripple solutions.
If $\tilde{b}=0$, necessarily $\tilde{c_{r}}=-v_{3,0,0}(K)$ by (79a). Then, we have
~~$\tilde{\tilde{\mathcal{B}}}(0,0,-v_{3,0,0}(K))=0,$ (81)
by construction. The corresponding Jacobian is
~~$\frac{\partial\tilde{\tilde{\mathcal{B}}}}{\partial(\tilde{b},\tilde{c_{r}})}(0,0,-v_{3,0,0}(K))=\begin{pmatrix}0&1\\ w_{2,1,0}(K)-v_{3,0,0}(K)&0\end{pmatrix},$ (82)
and a direct calculation confirms
$\textrm{det}\left(\begin{pmatrix}0&1\\ w_{2,1,0}(K)-v_{3,0,0}(K)&0\end{pmatrix}\right)=\frac{(K^{2}+1)(4K^{2}-61)}{6(K^{2}-4)(4K^{2}-1)}>0\quad\textrm{for}\quad K\geq 4.$ (83)
Thus, (82) is invertible, and the implicit function theorem gives us real analytic functions $\tilde{b}(a)$ and $\tilde{c_{r}}(a)$ such that ~~$\tilde{\tilde{\mathcal{B}}}(a,\tilde{b}(a),\tilde{c_{r}}(a))=0$, $\tilde{b}(0)=0$, and $\tilde{c_{r}}(0)=-v_{3,0,0}(K)$.
For each $K\geq 4$, we have a unique $a$-parameter family of solutions to (5) for sufficiently small $a\in\mathbb{R}$. These solutions travel at velocity
$c=c_{0}+a^{2}\tilde{c_{r}}(a)=\frac{K^{2}}{K^{2}+1}+a^{2}\tilde{c_{r}}(a),$ (84)
and take the form
$u(x;a)=a\Big{(}\cos(x)+\tilde{b}(a)\cos(Kx)\Big{)}+u_{r}\Big{(}x;a,a\tilde{b}(a),a^{2}\tilde{c_{r}}(a)\Big{)},$ (85)
where $u_{r}\in\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$ is real analytic in its parameters with Taylor expansion (35) and coefficients (92a)-(92c). The functions $\tilde{b}(a)$ and $\tilde{c_{r}}(a)$ are real analytic for sufficiently small $a$ and satisfy
$\tilde{b}(0)=0\quad\textrm{and}\quad\tilde{c_{r}}(0)=\frac{(K^{2}+1)(5K^{2}-24)}{6K^{2}(K^{2}-4)}.$ (86)
As before, one can show that $u(x;a)$ is real analytic in $a$ with values in $\textrm{H}^{\infty}_{\textrm{per},\textrm{even}}(\mathbb{T})$.
Unlike in previous cases, we cannot yet claim that (85) are Wilton ripple solutions. In particular, we must ensure that $\tilde{b}(a)$ does not identically vanish for all sufficiently small $a$. This is not so easy to verify. However, given we have rigorously proven that $u(x;a)$ and $c(a)$ are real analytic in $a$, one can return to (5), substitute in the appropriate expansions for $u(x;a)$ and $c(a)$, and apply a Poincaré-Linstedt perturbation method to obtain increasingly higher-order coefficients of these expansions. These calculations have been carried out in great detail by [14], where analyticity was assumed a priori. In Lemma 5 of this work, it was shown that $\mathcal{F}_{K}(u(x;a))\neq 0$ and, in fact, $\mathcal{F}_{K}(u(x;a))=\mathcal{O}(a^{K-2})$ as $a\rightarrow 0$. Hence, we conclude $\tilde{b}(a)\not\equiv 0$ and, more specifically, $\tilde{b}(a)=\mathcal{O}(a^{K-3})$ as $a\rightarrow 0$, completing the proof of the final case of Theorem 1.
3 Conclusions
Wilton ripples represent a special class of small-amplitude, periodic, traveling-wave solutions that bifurcate from two co-propagating cosine waves. In general, these solutions represent important resonances in the study of water waves [24] as well as in magnetohydrodynamics [26]. Current existence proofs of Wilton ripples [2]-[3], [12, 16], and [19]-[20] mostly consider the simplest resonant configuration, when the Wilton ripple bifurcates from a superposition of $\cos(x)$ and $\cos(2x)$. Among the primary challenges in proving the existence of the remaining Wilton ripples, which bifurcate from a superposition of $\cos(x)$ and $\cos(Kx)$ for any $K\in\mathbb{N}\setminus\{1\}$, is to prove their analyticity with respect to a small-amplitude parameter.
This paper develops a proof that shows for the first time the existence of all Wilton ripple solutions of a nonlinear dispersive equation, namely the Kawahara equation. The proof relies on a carefully constructed Lyapunov-Schmidt reduction to show the analyticity of each solution with respect to a small-amplitude parameter. High-order asymptotic expansions of these solutions, derived in [14], are then used to complete the proof. The author anticipates that this proof can be extended to show the existence of all Wilton ripples of many other nonlinear dispersive equations, including those for gravity-capillary water waves, provided the necessary asymptotic expansions are computed. It is also expected that this proof will connect the classic approach to Wilton ripples, where equation parameters like $\beta$ are fixed, and more contemporary approaches that study sheets of Wilton ripples [8], where the superposition coefficients $a$ and $b$ are instead fixed.
4 Appendix: The Taylor Coefficients $u_{r}^{(i,j,k)}(x)$
4.1 Case 1: $K=2$
We begin with the expansion of $u_{r}(x;a,b,c_{r})$ when $K=2$. Substituting (35) into (29) and equating collective powers in $a$, $b$, and $c_{r}$ yields a hierarchy of inhomogeneous differential equations that uniquely determines $u_{r}^{(i,j,k)}(x)$ for increasingly larger $i$, $j$, and $k$. Calculating these equations is straightforward but tedious, so we appeal to Mathematica for help. The interested reader can consult CompanionToKawaharaWiltonRipples.nb for further details.
The differential equations for $u_{r}^{(i,j,k)}(x)$ coincide when $i+j+k=1$ and are given by
$(c_{0}+\mathcal{L})u^{(i,j,k)}(x)=0.$ (87)
As $u^{(i,j,k)}(x)\in\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}$, we conclude
$u^{(i,j,k)}(x)=0\quad\textrm{for}\quad i+j+k=1.$ (88)
In other words, there are no first-order terms in the Taylor series expansion of $u_{r}(x;a,b,c_{r})$.
When $i+j+k=2$, we obtain six differential equations. In particular,
$(c_{0}+\mathcal{L})u^{(2,0,0)}(x)$ $=-\frac{1}{2},$ (89a)
$(c_{0}+\mathcal{L})u^{(1,1,0)}(x)$ $=-\cos(3x),$ (89b)
$(c_{0}+\mathcal{L})u^{(0,2,0)}(x)$ $=-\frac{1}{2}-\frac{1}{2}\cos(4x),$ (89c)
$(c_{0}+\mathcal{L})u^{(0,1,1)}(x)$ $=0,$ (89d)
$(c_{0}+\mathcal{L})u^{(0,0,2)}(x)$ $=0,$ (89e)
$(c_{0}+\mathcal{L})u^{(1,0,1)}(x)$ $=0.$ (89f)
By undetermined coefficients (or an equivalent method), one finds
$u^{(2,0,0)}(x)$ $=-\frac{1}{2c_{0}},$ (90a)
$u^{(1,1,0)}(x)$ $=-\frac{1}{c_{0}-9+81\beta}\cos(3x),$ (90b)
$u^{(0,2,0)}(x)$ $=-\frac{1}{2c_{0}}-\frac{1}{2\left(c_{0}-16+256\beta\right)}\cos(4x).$ (90c)
The remaining coefficient functions are zero by the same logic as in the $i+j+k=1$ case.
4.2 Case 2: $K\geq 3$
We proceed as in Case 1 but leave the details to Mathematica. When $i+j+k=1$, we find
$u^{(i,j,k)}(x)=0.$ (91)
When $i+j+k=2$, we have instead
$u^{(2,0,0)}(x)$ $=-\frac{1}{2c_{0}}-\frac{1}{2\left(c_{0}-4+16\beta\right)}\cos(2x)$ (92a)
$u^{(1,1,0)}(x)$ $=-\frac{1}{c_{0}-(K-1)^{2}+\beta(K-1)^{4}}\cos((K-1)x)-\frac{1}{c_{0}-(K+1)^{2}+\beta(K+1)^{4}}\cos((K+1)x),$ (92b)
$u^{(0,2,0)}(x)$ $=-\frac{1}{2c_{0}}-\frac{1}{2\left(c_{0}-4K^{2}+16\beta K^{4}\right)}\cos(2Kx),$ (92c)
and the remaining coefficient functions are zero.
Remark: If $K=2$, the coefficients of $\cos(2x)$ in (92a) and of $\cos((K-1)x)$ in (92b) are singular. This is to be expected, as $u^{(i,j,k)}(x)$ must be orthogonal to $\textrm{Null}\{c_{0}+\mathcal{L}\}=\textrm{Span}\{\cos(x),\cos(2x)\}$ when $K=2$. It is for this reason that this case is considered separately.